OpenStudy (anonymous):

This should be really easy, but I just dont get it! Please Help Simplify the expression. 18 + 24 ÷ (–4) • (–6) A. 54 B. –18 C. 19 D. 17 Thx!

OpenStudy (jdoe0001):

\(\bf \Large \text{PEMDAS}\) multiplication and division first, from left to right then additiona addition and subtraction, from left to right

OpenStudy (justuu):

A.54 I think

OpenStudy (anonymous):

I got 19 doing it from left to right, but how did u get 54??

OpenStudy (jdoe0001):

what did you get for 24 ÷ (–4) ?

OpenStudy (anonymous):

-6

OpenStudy (jdoe0001):

ok.. then.. one sec

OpenStudy (jdoe0001):

\(\large { \begin{array}{ccccccc} 18&+&24&\div&(-4)&\cdot &(-6)\\ &&&-6&&\cdot &-6\\ &&&&&+36\\ 18&&&&&+36\\ \\\hline\\ &&&54 \end{array} }\)